OpenStudy (anonymous):

Take r=4sin(theta) into rectangular (x,y) form

OpenStudy (anonymous):

\[\begin{cases}r=\sqrt{x^2+y^2}\\\theta=\tan^{-1}\dfrac{y}{x}\end{cases}\] \[\sqrt{x^2+y^2}=4\sin\left(\tan^{-1}\frac{y}{x}\right)\]  \[\sin\theta=\sin\left(\tan^{-1}\frac{y}{x}\right)=\frac{y}{\sqrt{x^2+y^2}}\] \[\begin{align*}\sqrt{x^2+y^2}&=\frac{4y}{\sqrt{x^2+y^2}}\\ x^2+y^2&=4y&\text{this form is fine, but if you want to continue...}\\ y^2-4y&=-x^2\\ y^2-4y+4&=4-x^2\\ (y-2)^2&=4-x^2\\ y&=2\pm\sqrt{4-x^2} \end{align*}\]